# Higher-Rank Irreducible Cartesian Tensors for Equivariant Message Passing

### v1

@@ B.3 Asymptotic computational complexity
We analyze the computational complexity of the irreducible Cartesian tensor product with respect to the maximum rank $L$ used. For each tuple of $\left(i_{1},\cdots,i_{L}\right)$ of a rank-$L$ tensor in Eq. (2), the single contraction term $\left(\mathbf{x}_{L}\cdot(k+m)\cdot\mathbf{y}_{L}\right)$ has a computational complexity of $\mathcal{O}\left(3^{k+m}\right)$. The total computational complexity is $\mathcal{O}\left(3^{L}\right)$ for even products and $\mathcal{O}\left(3^{L+1}\right)$ for odd ones—due to the additional double contraction with the Levi-Civita symbol. The computation of the set of permutations over the $L$ unsymmetrized indices scales as $L!/\left(2^{L/2}\left(L/2\right)!\right)$ for even products and $L!/\left(2^{\left(L-1\right)/2}\left(\left(L-1\right)/2\right)!\right)$ for odd ones [45]. Because each final rank-$L$ tensor has $3^{L}$ elements, the total complexity for computing an irreducible Cartesian tensor of rank $L$ is $\mathcal{O}\left(9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$.
@@
Recent state-of-the-art equivariant message-passing architectures employ tensors of rank $L\leq 4$ [33, 32, 78]. Also, physical properties are represented by tensors of rank $L\leq 4$. For these spherical tensors, the number of calculations required to obtain an element of a new $L$-rank tensor is $\left(2L+1\right)^{5}$, or $\left(2L+1\right)^{6}$ for a $\left(2L+1\right)$-dimensional object $\mathbf{Y}^{L}$. Thus, the irreducible Cartesian tensor product is expected to be more computationally efficient for $L\leq 4$, assuming similar multiplicative factors and negligible sub-leading terms. Compared to the Gaunt tensor product with $(2L+1)^{3}$, irreducible Cartesian tensors are expected to be less computationally efficient for $L\geq 3$. However, while the Gaunt tensor product is restricted only to even parity, i.e., it excludes tensors with odd $l_{1}+l_{2}-l_{3}=2k+1$ (e.g., pseudo-vectors), the irreducible Cartesian tensor product naturally allows their computation.
As regards the cost of performing message passing at each layer, the general neighbors’ aggregation of Eq. (4) has a cost of $\mathcal{O}\left(\mathcal{E}9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$, where $\mathcal{E}$ is the number of edges in the system. Similarly, computing the many-body features tensor via Eq. (6) requires to perform $\nu$ Cartesian tensor products for all $\mathcal{K}=\mathrm{len}(\eta_{\nu})$ possible $\nu$-fold tensor products, resulting in an asymptotic complexity of $\mathcal{O}\left(\mathcal{M}\mathcal{K}\nu 9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$ with $\mathcal{M}$ denoting the number of nodes. Thus, the computational complexity for one ICTP message-passing layer is $\mathcal{O}\left(\left(\mathcal{E}+\mathcal{M}\mathcal{K}\nu\right)9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$.
@@
Finally, memory consumption for Cartesian tensors is often believed to be less advantageous than that of spherical tensors. However, for contracting two rank-$L$ spherical tensors, space complexity is about $(2L+1)^{2}$, resulting from the definition of the Clebsch–Gordan tensor product in Section 2, i.e., due to the computation of intermediate tensors $\mathbf{Y}^{L}\otimes\mathbf{Y}^{L}\in\mathbb{R}^{(2L+1)\times(2L+1)}$. For a Cartesian tensor, the space complexity is about $3^{L}$. Thus, the irreducible Cartesian tensor product can be expected to require the same or less memory for $L\leq 4$ than the Clebsch–Gordan counterpart.
@@ Appendix C Proof of Cartesian message-passing equivariance to actions of the orthogonal group
$\begin{split}&\left(\left(R\mathbf{T}_{l_{1}}\right)\otimes\left(R\mathbf{T}_{l_{2}}\right)\right)_{i_{1}\cdots i_{l_{1}}i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}=\\ &\left(R\mathbf{T}_{l_{1}}\right)_{i_{1}\cdots i_{l_{1}}}\left(R\mathbf{T}_{l_{2}}\right)_{i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}=\\ &\left(\sum_{j_{1}}\cdots\sum_{j_{l_{1}}}R_{i_{1}j_{1}}\cdots R_{i_{l}j_{l_{1}}}\left(\mathbf{T}_{l_{1}}\right)_{j_{1}\cdots j_{l_{1}}}\right)\left(\sum_{j_{l_{1}+1}}\cdots\sum_{j_{l_{1}+l_{2}}}R_{i_{l_{1}+1}j_{l_{1}+1}}\cdots R_{i_{l_{1}+l_{2}}j_{l_{1}+l_{2}}}\left(\mathbf{T}_{l_{2}}\right)_{j_{l_{1}+1}\cdots j_{l_{1}+l_{2}}}\right)=\\ &\sum_{j_{1}}\cdots\sum_{j_{l_{1}}}\sum_{j_{l_{1}+1}}\cdots\sum_{j_{l_{1}+l_{2}}}R_{i_{1}j_{1}}\cdots R_{i_{l_{1}}j_{l_{1}}}R_{i_{l_{1}+1}j_{l_{1}+1}}\cdots R_{i_{l_{1}+l_{2}}j_{l_{1}+l_{2}}}\left(\mathbf{T}_{l_{1}}\right)_{j_{1}\cdots j_{l_{1}}}\left(\mathbf{T}_{l_{2}}\right)_{j_{l_{1}+1}\cdots j_{l_{1}+l_{2}}}\\ &=\left(R\left(\mathbf{T}_{l_{1}}\otimes\mathbf{T}_{l_{2}}\right)\right)_{i_{1}\cdots i_{l_{1}}i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}.\end{split}$
@@
$\begin{split}&\left(\left(R\mathbf{x}_{l_{1}}\right)\cdot(s)\cdot\left(R\mathbf{y}_{l_{2}}\right)\right)_{\beta_{1}\cdots\beta_{l_{1}-s}\delta_{1}\cdots\delta_{l_{2}-s}}=\\ &\sum_{\alpha_{1},\cdots,\alpha_{s}}\left(R\mathbf{x}_{l_{1}}\right)_{\alpha_{1}\cdots\alpha_{s}\beta_{1}\cdots\beta_{l_{1}-s}}\left(R\mathbf{y}_{l_{2}}\right)_{\alpha_{1}\cdots\alpha_{s}\delta_{1}\cdots\delta_{l_{2}-s}}=\\ &\sum_{\alpha_{1},\cdots,\alpha_{s}}\left(\sum_{\begin{subarray}{c}\gamma_{1},\cdots,\gamma_{s}\\ \eta_{1},\cdots,\eta_{l_{1}-s}\end{subarray}}R_{\alpha_{1}\gamma_{1}}\cdots R_{\alpha_{s}\gamma_{s}}R_{\beta_{1}\eta_{1}}\cdots R_{\beta_{l_{1}-s}\eta_{l_{1}-s}}\left(\mathbf{x}_{l_{1}}\right)_{\gamma_{1}\cdots\gamma_{s}\eta_{1}\cdots\eta_{l_{1}-s}}\right)\times\\ &\hskip 34.1433pt\left(\sum_{\begin{subarray}{c}\tilde{\gamma}_{1},\cdots,\widetilde{\gamma}_{s}\\ \widetilde{\eta}_{1},\cdots,\widetilde{\eta}_{l_{2}-s}\end{subarray}}R_{\alpha_{1}\widetilde{\gamma}_{1}}\cdots R_{\alpha_{s}\widetilde{\gamma}_{s}}R_{\delta_{1}\widetilde{\eta}_{1}}\cdots R_{\delta_{l_{2}-s}\widetilde{\eta}_{l_{2}-s}}\left(\mathbf{y}_{l_{2}}\right)_{\widetilde{\gamma}_{1}\cdots\widetilde{\gamma}_{s}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}}\right)=\\ &\left(\sum_{\begin{subarray}{c}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{1}-s}\end{subarray}}R_{\beta_{1}\eta_{1}}\cdots R_{\beta_{l_{1}-s}\eta_{l_{1}-s}}\right)\left(\sum_{\begin{subarray}{c}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}\end{subarray}}R_{\delta_{1}\widetilde{\eta}_{1}}\cdots R_{\delta_{l_{2}-s}\widetilde{\eta}_{l_{2}-s}}\right)\times\\ &\hskip 135.15059pt\sum_{\begin{subarray}{c}\gamma_{1}\cdots\gamma_{s}\end{subarray}}\left(\mathbf{x}_{l_{1}}\right)_{\gamma_{1}\cdots\gamma_{s}\eta_{1}\cdots\eta_{l_{1}-s}}\left(\mathbf{y}_{l_{2}}\right)_{\gamma_{1}\cdots\gamma_{s}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}}=\\ &\left(R\left(\mathbf{x}_{l_{1}}\cdot(s)\cdot\mathbf{y}_{l_{2}}\right)\right)_{\beta_{1}\cdots\beta_{l_{1}-s}\delta_{1}\cdots\delta_{l_{2}-s}}.\end{split}$
@@
      $\big{(}\mathbf{A}_{ukl_{1}}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}=\sum_{v\in\mathcal{N}(u)}R_{kl_{1}}^{(1)}(r_{uv})\big{(}\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}W_{kZ_{v}}$
@@
    $\big{(}\mathbf{A}_{ukl_{3}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{3}}}=\sum_{v\in\mathcal{N}(u)}\Big{(}R_{kl_{1}l_{2}l_{3}}^{(t)}(r_{uv})\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\otimes_{\mathrm{Cart}}\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}l_{2}}^{(t)}\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}\Big{)}_{i_{1}i_{2}\cdots i_{l_{3}}},$

### v2

@@ B.3 Computational complexity and runtime analysis
We analyze the computational complexity of the irreducible Cartesian tensor product with respect to the maximum rank $L$ used. For each tuple of $\left(i_{1},\cdots,i_{L}\right)$ of a rank-$L$ tensor in Eq. (2), the single contraction term $\left(\mathbf{x}_{L}\cdot(k+m)\cdot\mathbf{y}_{L}\right)$ has a cost of $3^{k+m}$. The total computational cost is $3^{L}$ for even tensor products and $3^{L+1}$ for odd ones—due to the additional double contraction with the Levi-Civita symbol. The computation of the set of permutations over the $L$ unsymmetrized indices scales as $L!/\left(2^{L/2}\left(L/2\right)!\right)$ for even tensor products and $L!/\left(2^{\left(L-1\right)/2}\left(\left(L-1\right)/2\right)!\right)$ for odd ones [45]. Because each final rank-$L$ tensor has $3^{L}$ elements, the complexity for computing an irreducible Cartesian tensor of rank $L$ is $\mathcal{O}\left(9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$. This expression is used throughout this work as it captures contributions from tensor contractions and index permutations, though it simplifies asymptotically to $\mathcal{O}\left(L^{L}\right)$ using Stirling’s approximation for factorials.
@@
The number of calculations required to obtain a rank-$L$ spherical tensor through the Clebsch–Gordan tensor product is $\left(2L+1\right)^{5}$. Thus, the computational complexity of the Clebsch–Gordan tensor product is $\mathcal{O}\left(L^{5}\right)$. While the Clebsch–Gordan tensor product is more computationally efficient than the irreducible Cartesian tensor product for $L\rightarrow\infty$, the latter is expected to be advantageous for smaller tensor ranks, i.e., $L\leq 4$, assuming similar multiplicative factors and negligible sub-leading terms. Tensors of rank $L\leq 4$ are particularly relevant for equivariant message-passing architectures and representing physical properties [33, 32, 72]. Compared to the Gaunt tensor product with the computational complexity of $\mathcal{O}\left(L^{3}\right)$, including all $\left(l_{1},l_{2}\right)\rightarrow l_{3}$ [36], the irreducible Cartesian tensor product may be less advantageous for $L\geq 3$. However, the Gaunt-coefficients-based approach excludes odd tensor products, limiting the expressive power and the range of possible architectures.
As regards the cost of performing message passing at each layer, the general neighbors’ aggregation of Eq. (4) has a cost of $\mathcal{E}N_{\mathrm{ch}}9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)$, where $\mathcal{E}$ is the number of edges in the atomic system and $N_{\mathrm{ch}}$ is the number of feature channels. Computing many-body features via Eq. (6) requires to iteratively perform $\nu-1$ Cartesian tensor products for all $\mathcal{K}=\mathrm{len}(\eta_{\nu})$ possible $\nu$-fold tensor products, resulting in a computational cost of $\mathcal{M}N_{\mathrm{ch}}\mathcal{K}(9^{L}L!/(2^{L/2}(L/2)!))^{\nu-1}$ with $\mathcal{M}$ denoting the number of nodes. For the Clebsch–Gordan tensor product, the corresponding number of calculations is $\mathcal{E}N_{\text{ch}}L^{5}$ and $\mathcal{M}N_{\text{ch}}\mathcal{K}L^{5(\nu-1)}$. Spherical models, however, can use generalized Clebsch–Gordan coefficients, resulting in $\mathcal{M}N_{\text{ch}}\mathcal{K}L^{\frac{1}{2}\nu(\nu+3)}$ for the product basis. We can remove the factor $\mathcal{K}$ from the above expression by restricting the parameterization to uncoupled feature channels as in Eq. (A1) and obtain $\mathcal{M}N_{\text{ch}}L^{\frac{1}{2}\nu(\nu+3)}$. Thus, MACE with this specific choice of the product basis can be more computationally efficient than ICTP only for large $N_{\text{ch}}$ and small $\nu$, given $L\leq 4$. In this work, we have shown that leveraging the symmetry of the irreducible Cartesian tensor product and coupled feature channels can improve the computational cost of ICTP. The former, in particular, reduces the effective number of $\nu$-fold tensor products $\mathcal{K}$. Further optimization of the architecture, however, is possible and is expected to fully exploit the advantage of Cartesian tensors.
@@
Finally, memory consumption for Cartesian tensors is often believed to be less advantageous than that of spherical tensors. However, for contracting two rank-$L$ spherical tensors, the memory requirement is about $(2L+1)^{2}$, resulting from the definition of the Clebsch–Gordan tensor product in Section 2. For a Cartesian tensor, the memory requirement is about $3^{L}$. Thus, the irreducible Cartesian tensor product can be expected to require the same or less memory for $L\leq 4$ than the Clebsch–Gordan counterpart. The memory consumption of models based on spherical tensors is further increased by the use of generalized Clebsch–Gordan coefficients, as we demonstrate in Section 5.
@@ Appendix C Proof of Cartesian message-passing equivariance to actions of the orthogonal group
$\begin{split}&\left(\left(R\mathbf{T}_{l_{1}}\right)\otimes\left(R\mathbf{T}_{l_{2}}\right)\right)_{i_{1}\cdots i_{l_{1}}i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}\\ &\quad\quad=\left(R\mathbf{T}_{l_{1}}\right)_{i_{1}\cdots i_{l_{1}}}\left(R\mathbf{T}_{l_{2}}\right)_{i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}\\ &\quad\quad=\left(\sum_{j_{1},\dots,j_{l_{1}}}R_{i_{1}j_{1}}\cdots R_{i_{l}j_{l_{1}}}\left(\mathbf{T}_{l_{1}}\right)_{j_{1}\cdots j_{l_{1}}}\right)\left(\sum_{j_{l_{1}+1},\dots,j_{l_{1}+l_{2}}}R_{i_{l_{1}+1}j_{l_{1}+1}}\cdots R_{i_{l_{1}+l_{2}}j_{l_{1}+l_{2}}}\left(\mathbf{T}_{l_{2}}\right)_{j_{l_{1}+1}\cdots j_{l_{1}+l_{2}}}\right)\\ &\quad\quad=\sum_{j_{1}}\cdots\sum_{j_{l_{1}}}\sum_{j_{l_{1}+1}}\cdots\sum_{j_{l_{1}+l_{2}}}R_{i_{1}j_{1}}\cdots R_{i_{l_{1}}j_{l_{1}}}R_{i_{l_{1}+1}j_{l_{1}+1}}\cdots R_{i_{l_{1}+l_{2}}j_{l_{1}+l_{2}}}\left(\mathbf{T}_{l_{1}}\right)_{j_{1}\cdots j_{l_{1}}}\left(\mathbf{T}_{l_{2}}\right)_{j_{l_{1}+1}\cdots j_{l_{1}+l_{2}}}\\ &\quad\quad=\left(R\left(\mathbf{T}_{l_{1}}\otimes\mathbf{T}_{l_{2}}\right)\right)_{i_{1}\cdots i_{l_{1}}i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}.\end{split}$
@@
$\begin{split}&\left(\left(R\mathbf{x}_{l_{1}}\right)\cdot(s)\cdot\left(R\mathbf{y}_{l_{2}}\right)\right)_{\beta_{1}\cdots\beta_{l_{1}-s}\delta_{1}\cdots\delta_{l_{2}-s}}\\ &\quad\quad=\sum_{\alpha_{1},\cdots,\alpha_{s}}\left(R\mathbf{x}_{l_{1}}\right)_{\alpha_{1}\cdots\alpha_{s}\beta_{1}\cdots\beta_{l_{1}-s}}\left(R\mathbf{y}_{l_{2}}\right)_{\alpha_{1}\cdots\alpha_{s}\delta_{1}\cdots\delta_{l_{2}-s}}\\ &\quad\quad=\sum_{\alpha_{1},\cdots,\alpha_{s}}\left(\sum_{\begin{subarray}{c}\gamma_{1},\cdots,\gamma_{s}\\ \eta_{1},\cdots,\eta_{l_{1}-s}\end{subarray}}R_{\alpha_{1}\gamma_{1}}\cdots R_{\alpha_{s}\gamma_{s}}R_{\beta_{1}\eta_{1}}\cdots R_{\beta_{l_{1}-s}\eta_{l_{1}-s}}\left(\mathbf{x}_{l_{1}}\right)_{\gamma_{1}\cdots\gamma_{s}\eta_{1}\cdots\eta_{l_{1}-s}}\right)\\ &\hskip 66.86397pt\times\left(\sum_{\begin{subarray}{c}\tilde{\gamma}_{1},\cdots,\widetilde{\gamma}_{s}\\ \widetilde{\eta}_{1},\cdots,\widetilde{\eta}_{l_{2}-s}\end{subarray}}R_{\alpha_{1}\widetilde{\gamma}_{1}}\cdots R_{\alpha_{s}\widetilde{\gamma}_{s}}R_{\delta_{1}\widetilde{\eta}_{1}}\cdots R_{\delta_{l_{2}-s}\widetilde{\eta}_{l_{2}-s}}\left(\mathbf{y}_{l_{2}}\right)_{\widetilde{\gamma}_{1}\cdots\widetilde{\gamma}_{s}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}}\right)\\ &\quad\quad=\left(\sum_{\begin{subarray}{c}\eta_{1}\cdots\eta_{l_{1}-s}\end{subarray}}R_{\beta_{1}\eta_{1}}\cdots R_{\beta_{l_{1}-s}\eta_{l_{1}-s}}\right)\left(\sum_{\begin{subarray}{c}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}\end{subarray}}R_{\delta_{1}\widetilde{\eta}_{1}}\cdots R_{\delta_{l_{2}-s}\widetilde{\eta}_{l_{2}-s}}\right)\\ &\hskip 66.86397pt\times\sum_{\begin{subarray}{c}\gamma_{1}\cdots\gamma_{s}\end{subarray}}\left(\mathbf{x}_{l_{1}}\right)_{\gamma_{1}\cdots\gamma_{s}\eta_{1}\cdots\eta_{l_{1}-s}}\left(\mathbf{y}_{l_{2}}\right)_{\gamma_{1}\cdots\gamma_{s}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}}\\ &\quad\quad=\left(R\left(\mathbf{x}_{l_{1}}\cdot(s)\cdot\mathbf{y}_{l_{2}}\right)\right)_{\beta_{1}\cdots\beta_{l_{1}-s}\delta_{1}\cdots\delta_{l_{2}-s}}.\end{split}$
@@
      $\big{(}\mathbf{A}_{ukl_{1}}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}=\sum\nolimits_{v\in\mathcal{N}(u)}R_{kl_{1}}^{(1)}(r_{uv})\big{(}\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}W_{kZ_{v}}$
@@
    $\big{(}\mathbf{A}_{ukl_{3}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{3}}}=\sum\nolimits_{v\in\mathcal{N}(u)}\Big{(}R_{kl_{1}l_{2}l_{3}}^{(t)}(r_{uv})\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\otimes_{\mathrm{Cart}}\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}l_{2}}^{(t)}\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}\Big{)}_{i_{1}i_{2}\cdots i_{l_{3}}},$